3. Show that if n e Z so that n is odd then 8|n2 + (n + 6)2 +6. 4. (a) Let a, b, and n be integers so that n > 2. Define: а iPlease answer all!!
Answer #1

sol: Any odd integer n is of the form I n=2k+I , where KEEL Now n²+ (n+6² +6 = (2K+ 1)²+(2k +1+6)² +6 {: n=2k+1} = 4K?+ 46+1+(dNo, 27€ 19 (mod 8). Explanation. Since 24-13 = 14 which is not divisible by 8 So 27 is not congruent to 13 modulo 8. We wri